OpenStudy (anonymous):

If h(x) = −4x − 10, find h(−5).

OpenStudy (anonymous):

plug -5 in x

OpenStudy (anonymous):

Like this -4(-5) - 10 then you will have your answer

OpenStudy (anonymous):

You simply have to substitute the value of x as - 5 in the function and solve for x

OpenStudy (anonymous):

h(x) = −4x − 10 So h(-5)= -4(-5)-10 = 20-10 = 10